Denise has utility over consumption c and leisure l defined by the following function:

U(c, l) = c + l

a) Suppose Denise has two units of consumption and three units of leisure. What is her utility?

b) Suppose Denise has four units of consumption and one unit of leisure. What is her utility?

c) Graph her indifference curves. Draw at least three separate indifference curves, for U = {2, 4, 6}. Label your axes accordingly.
Answer #1

a) Utility at C = 2 and L = 3 is U = 2 + 3 = 5 units

b) Utility at C = 4 and L = 1 is U = 4 + 1 = 5 units

c) Since consumption and leisure are two goods with perfect substitution the indifference curves are straight lines.